A 22.8 kg box is placed at the bottom of a 39.7° ramp whose surface is rough enough that the coefficient of kinetic friction is 0.222. You push on the box parallel to the ramp's surface, so that the box accelerates up the ramp at 0.911 m/s2. What must be the force (in N) supplied by you?
Answer #1

considering equilibrium of forces along the ramp

ma = F - mg sin x - u mg cos x

22.8 * 0.911 = F - 22.8 * 9.8* sin 39.7 - 0.222* 22.8* 9.8* cos 39.7

F = 201.662 N

========

Comment before rate in case any doubt, will reply for sure.. Goodluck